OpenStudy (anonymous):

PLEASE HELP ME SOLVE X+484% OF X=14.6 BILLION

OpenStudy (amistre64):

distrbute the x out, and divide off the excess

OpenStudy (anonymous):

x + 484% of x = 14.6 billion (" of " means multiply) x + 4.84x = 14.6 billion (combine like terms) 5.84x = 14.6 billion x = 14.6/5.84 x = 2.5 billion check... x + 4.84x = 14.6 billion 2.5 + (4.84 * 2.5) = 14.6 2.5 + 12.1 = 14.6 14.6 billion = 14.6 billion (correct) If I did this correctly....the answer is 2.5 billion ?